OpenStudy (anonymous):

The average time it takes to manufacture a car in the United States is equal to the average time it takes to manufacture a car in Japan plus 8.1 hours. The average time it takes to manufacture a car in the United States is 24.9 hours

OpenStudy (anonymous):

I'm kind of out of it today. Don't know if it is l+8.1=24.9 or l-8.1=24.9...

OpenStudy (anonymous):

there's no question here, just a collection of statements lol

OpenStudy (anonymous):

Oh, I want to know how to write in an expression...

OpenStudy (anonymous):

oh, ok: Let U=average USA time Let J=average Japan time then, U=J+8.1 but U is given as U=24.9 hours So, 24.9=J+8.1 J=16.8 hours